# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ A.3 Non-uniform convergence rates
Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$. Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $X$ is a martingale that is bounded from below by $x$. By Ville’s inequality,
@@
$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{-k}.$
@@
Recall from Appendix A.2 that the rates obtained in Theorem 16 are asymptotically optimal for the Simons and median martingales (see Examples 17 and 18). For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
@@
Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $\sup_{A}f\leqslant(1+\varepsilon)\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit). Then for any interval $a,b]\subseteq[0,1]$,
$\mathbb{E}[U\mid U\in[a,b]]\in\Big{[}a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)\Big{]}.$
@@
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\mathrm{d}x$
$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
@@
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])\geqslant\inf_{m\in[a+\frac{1}{2+\varepsilon}(b-a),\,a+\frac{1+\varepsilon}{2+\varepsilon}(b-a)]}\int_{a}^{b}(x-m)^{2}\mathrm{d}x\geqslant\frac{1}{12}(b-a)^{3}.$
@@
$\frac{1}{12}x_{*}^{3}\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(1-x_{*})^{3}\quad\text{and}\quad\frac{1}{12}(1-x_{*})^{3}\geqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}x_{*}^{3}.$
@@
A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
$\mathbb{E}[(U-M_{1})^{2}]\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(x_{*}^{3}+(1-x_{*})^{3})\leqslant\frac{4(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(\frac{1}{2}+a_{\varepsilon})^{3}.$
@@
$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{2(1+\varepsilon)^{4}}{(2+\varepsilon)^{3}}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
@@ A.4 Proof of Theorem 16
Let $Q$ be the quantile function of $U$. Since $U$ is atomless, we have if $u=Q(p)$,
@@
Since $U$ is atomless, it is non-degenerate, and hence $m_{b}(p_{\mathrm{var}})-m_{a}(p_{\mathrm{var}})>0$ and $p_{\mathrm{var}}\in(0,1)$. Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but have limits on both sides at every point. Since $p_{\mathrm{var}}\in\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p)$, there are two cases:
@@
If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$.
@@
Recall that our goal is to bound (37). The probability of the mass at each step is multiplied by $q$, and the edge lengths are bounded by $(a+b)/2$ and $(a^{\prime}+b^{\prime})/2$. By (39), their product multiplies by at most $1/2$ each time we update the tuple. We have thus shown that
$\max_{1\leqslant j\leqslant 2^{k-1}}(p_{2j-1}d_{2j-1}^{2}+p_{2j}d_{2j}^{2})\leqslant 2^{1-k}.$ (40)
Observe also that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$ (because the edges at a single level cannot intersect each other since they represent the conditional means of $U$ in disjoint intervals). By Hölder’s inequality,
$\begin{split}\mathbb{E}[(M_{k-1}-M_{k})^{2}]=\sum_{j=1}^{2^{k}}p_{j}d_{j}^{2}&\leqslant\Big{(}\sum_{j=1}^{2^{k}}p_{j}d_{j}^{3}\Big{)}^{2/3}\Big{(}\sum_{j=1}^{2^{k}}p_{j}\Big{)}^{1/3}\\ &\leqslant\Big{(}\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})\sum_{j=1}^{2^{k}}d_{j}\Big{)}^{2/3}\leqslant\max_{j\in[2^{k}]}(p_{j}d_{j}^{2})^{2/3}\leqslant 2^{-2(k-1)/3}.\end{split}$ (41)
By the martingale property, we then have
$\mathbb{E}[(U-M_{k})^{2}]=\sum_{j=k}^{\infty}\mathbb{E}[(M_{j}-M_{j+1})^{2}]\leqslant\sum_{j=k}^{\infty}2^{-2j/3}\leqslant\frac{2^{-2k/3}}{1-2^{-2/3}}\leqslant 2.71\cdot 2^{-2k/3},$
as desired. ∎
@@
It is noted in Zhang et al. (2024) that the Simons martingale satisfies the separated tree condition (i.e., the branches of the tree representation from all different levels, when projected onto the real line as intervals, are either disjoint or have a containment relationship), which is a consequence of the construction. We follow a similar argument to the variance martingale. Consider a node $\mathbf{M}$ at level k1k-1 and a path $(\emptyset=\mathbf{N}_{0},\dots,\mathbf{N}_{k-1}=\mathbf{M})$ leading to $\mathbf{M}$, where $\emptyset$ denotes the root of a binary tree. We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\leavevmode\nobreak\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$. The quantities $A_{j}$ and $B_{j}$ represent the maximum possible lengths of the two edges emanating from a node $\mathbf{N}_{j}$, and $q$ represents the one-step splitting probability leading to $\mathbf{N}_{j}$, based on the constraints from the separated tree property.
We start from $(A_{1},B_{1},q_{1})=(1,1,1)$. Again, denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants at level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. Consider the node $\mathbf{N}_{j}$ at level $j$. Then $(A_{j},B_{j},q_{j})$ must be of the form (while not distinguishing the order of $A_{j}$ and $B_{j}$)
@@
$A_{k}B_{k}\prod_{\ell=1}^{k-1}q_{\ell}=\prod_{\ell=1}^{k-1}\frac{A_{\ell}B_{\ell}}{A_{\ell-1}B_{\ell-1}}\prod_{\ell=1}^{k-1}q_{\ell}\leqslant\prod_{\ell=1}^{k-1}\Big{(}\frac{1}{q_{\ell}}-1\Big{)}q_{\ell}=\prod_{\ell=1}^{k-1}(1-q_{\ell}),$
@@
By the martingale property,
@@
At level $k$, there are $2^{k}$ vertices (denoted by $j\in[2^{k}]$) in the binary tree representation of the partition-based martingale approximation. Denote the probabilities of the vertices by $p_{j}$ (which correspond to $\mathbb{P}(U\in A_{j}),\leavevmode\nobreak\ A_{j}\in\pi_{k}$) and the locations by $\ell_{j}$. It follows that
@@
Note that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$. It follows analogously as the Hölder’s inequality argument in (41) that
@@ B.1 Proofs of results in Section 2
Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$, it holds that
@@
Let $C=(1+\delta)^{3}/\delta^{3}$. Without loss of generality, we assume that $\mathbb{E}[Y]=0$. Equivalent to (43) is
@@
To this end, we need an upper bound of $\mathrm{Var}(Y)$. Fix $g\in\mathcal{G}$ and let $g_{*}:=(\sup g+\inf g)/2$. By Minkowski’s inequality,
$\begin{split}\mathrm{Var}(Y)^{1/2}=\min_{y\in\mathbb{R}}\mathbb{E}[(Y-y)^{2}]^{1/2}\leqslant\mathbb{E}[(Y-g_{*})^{2}]^{1/2}&\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\mathbb{E}[(g(\mathbf{X})-g_{*})^{2}]^{1/2}\\ &\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\left\lVert g\right\rVert_{\mathrm{TV}}.\end{split}$ (46)
@@
It remains to establish (45). Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ by the cyclic minimax construction. Note that by construction, $J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the box $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$. By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$. Define $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
@@
    $y_{k,j}\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
    $V_{j}=\min\big{\{}(1+\delta^{-1})\,\mathbb{E}[(g(\mathbf{X})-y_{k,j})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\big{\}}.$
@@
    $y_{k,j}\not\in\mathrm{Ran}(g_{j})$. Define $U_{j}=(1+\delta)\,\mathbb{E}[(Y-g(\mathbf{X}))^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]$ and
    $V_{j}=\min\Big{\{}\frac{(1+\delta)^{3}}{\delta^{3}}\,p_{j}(\Delta g_{j})^{2},\,\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}]\Big{\}}.$
@@
as desired. On the other hand, suppose that $k=dq+r$ for some integer $q$ and $0\leqslant r<d$, and denote by $\pi_{k}$ the resulting partition of $\mathbb{R}^{d}$. Observe (by induction on $q,r$, and elementary geometry) that
@@
Recall that $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. The only difference from the proof of Theorem 2 is (52), since for non-marginally atomless measures (10) and (11) may not hold. Instead, by our assumptions, (10) is be replaced by
$\max\Big{\{}\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big{\}}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N},$
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$. As a consequence, if we denote by
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$

### v3

@@ A.3 Non-uniform convergence rates
Since the supports of $\{M_{k}\}_{k\geqslant 0}$ correspond to the means of $U$ conditioned on nested partitions of $\mathbb{R}$, we have $\tau_{M}^{+}<\infty$ implies $M_{k}\geqslant x$ for all $k\geqslant 0$ and $M_{1}=y$. Therefore, conditional on the event $\tau_{M}^{+}<\infty$, $\{M_{k}\}_{k\geqslant 0}$ is a martingale that is bounded from below by $x$. By Ville’s inequality,
@@
$\mathbb{E}[(U-M_{k})^{2}]\leqslant C(M^{-\delta^{\prime}}+M^{2}r^{k})\leqslant Cq^{k}.$
@@
Recall from Appendix A.2 that the rates obtained in Theorem 18 are asymptotically optimal for the Simons and median martingales (see Examples 19 and 20). For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
@@
Since one-step variance is always better than one-step minimax, it remains to show that for an interval $A$ with $(1-\varepsilon^{2})\sup_{A}f\leqslant\inf_{A}f$, there exists $C(\varepsilon)\downarrow 1/4$ as $\varepsilon>0$ such that at each step the total variance reduces by a factor of $C(\varepsilon)$ for the minimax martingale.
To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1-\varepsilon,1+\varepsilon]$, where $\varepsilon\in(0,1)$. Then for any interval $a,b]\subseteq[0,1]$,
$\mathbb{E}[U\mid U\in[a,b]]\in\Big[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)\Big].$
@@
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\leqslant\sup_{m\in[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)]}\int_{a}^{b}(1+\varepsilon)(x-m)^{2}\,\mathrm{d}x$
$\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}(b-a)^{3}.$
@@
$\mathbb{P}(U\in[a,b])\mathrm{Var}(U\mid U\in[a,b])$ $\geqslant\inf_{m\in[a+\frac{1-\varepsilon}{2}(b-a),\,a+\frac{1+\varepsilon}{2}(b-a)]}\int_{a}^{b}(1-\varepsilon)(x-m)^{2}\,\mathrm{d}x$
$\geqslant\frac{1-\varepsilon}{12}(b-a)^{3}.$
@@
$\frac{1-\varepsilon}{12}x_{*}^{3}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}(1-x_{*})^{3}\quad\text{and}\quad\frac{1-\varepsilon}{12}(1-x_{*})^{3}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}x_{*}^{3}.$
@@
A direct computation yields $\mathrm{Var}(U)\geqslant\frac{1-\varepsilon}{12}$. The total remaining variance $\mathbb{E}[(U-M_{1})^{2}]$ is bounded by
$\mathbb{E}[(U-M_{1})^{2}]$ $\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{12}\bigl(x_{*}^{3}+(1-x_{*})^{3}\bigr)$
$\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{6}\Big(\frac{1}{2}+a_{\varepsilon}\Big)^{3}.$
@@
$\frac{\mathbb{E}[(U-M_{1})^{2}]}{\mathrm{Var}(U)}\leqslant\frac{(1+\varepsilon)(1+3\varepsilon^{2})}{4(1-\varepsilon)}(1+2a_{\varepsilon})^{3}\to\frac{1}{4},$
@@ A.4 Proof of Theorem 18
Let $Q$ be the quantile function of $U$. Since $U$ is atomless, if $u=Q(p)$, then
@@
Since $U$ is atomless, it is non-degenerate, and hence $m_{b}(p_{\mathrm{var}})-m_{a}(p_{\mathrm{var}})>0$ and $p_{\mathrm{var}}\in(0,1)$. Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but has limits on both sides at every point. Since $p_{\mathrm{var}}\in\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p)$, there are two cases:
@@
If the support of $U$ is a connected interval, Lemma A.1 is a special case of Theorem 1 of Ishwaran (2015) applied with $f$ monotone and $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$.
@@
For each interval $I\in\pi_{k}$, write
$p_{I}:=\mathbb{P}(U\in I),\qquad m_{I}:=\mathbb{E}[U\mid U\in I],\qquad L_{I}:=m_{I}-\inf I,\qquad R_{I}:=\sup I-m_{I}.$
The same recursive argument leading to (40) yields
$\max_{I\in\pi_{k}}p_{I}L_{I}R_{I}\leqslant 2^{-2-k}.$ (41)
Indeed, along each update, the quantity $p_{I}L_{I}R_{I}$ is multiplied by at most $1/2$, while initially $L_{I}+R_{I}\leqslant\sup U-\inf U\leqslant 1$, so $L_{I}R_{I}\leqslant 1/4$.
Since $U\mid(U\in I)$ is supported on the interval $I$, we have
$\mathrm{Var}(U\mid U\in I)\leqslant(m_{I}-\inf I)(\sup I-m_{I})=L_{I}R_{I}.$
Therefore,
$\mathbb{E}[(U-M_{k})^{2}]$ $=\sum_{I\in\pi_{k}}p_{I}\mathrm{Var}(U\mid U\in I)\leqslant\sum_{I\in\pi_{k}}p_{I}L_{I}R_{I}.$
Set $s_{I}:=\sqrt{L_{I}R_{I}}$. By Hölder’s inequality and since $\sum_{I\in\pi_{k}}p_{I}=1$,
$\sum_{I\in\pi_{k}}p_{I}L_{I}R_{I}=\sum_{I\in\pi_{k}}p_{I}s_{I}^{2}$ $\leqslant\Big(\sum_{I\in\pi_{k}}p_{I}s_{I}^{3}\Big)^{2/3}\Big(\sum_{I\in\pi_{k}}p_{I}\Big)^{1/3}$
$\leqslant\Big(\max_{I\in\pi_{k}}p_{I}s_{I}^{2}\sum_{I\in\pi_{k}}s_{I}\Big)^{2/3}.$
Moreover,
$2s_{I}\leqslant L_{I}+R_{I}=\sup I-\inf I,$
so, since the intervals in $\pi_{k}$ are disjoint and contained in $\inf U,\sup U]$,
$\sum_{I\in\pi_{k}}s_{I}\leqslant\frac{1}{2}\sum_{I\in\pi_{k}}(\sup I-\inf I)\leqslant\frac{1}{2}(\sup U-\inf U)\leqslant\frac{1}{2}.$
Combining this with (41), we obtain
$\mathbb{E}[(U-M_{k})^{2}]\leqslant\Big(2^{-2-k}\cdot\frac{1}{2}\Big)^{2/3}\leqslant 2^{-2-2k/3},$
and the proof is complete. ∎
@@
It is noted in Zhang et al. (2024) that the Simons martingale satisfies the separated tree condition (i.e., the branches of the tree representation from all different levels, when projected onto the real line as intervals, are either disjoint or have a containment relationship), which is a consequence of the construction. We follow a similar argument to the variance martingale. Consider a node $\mathbf{M}$ at level k1k-1 and a path $(\emptyset=\mathbf{N}_{0},\dots,\mathbf{N}_{k-1}=\mathbf{M})$ leading to $\mathbf{M}$, where $\emptyset$ denotes the root of a binary tree. We provide an algorithm that recursively defines the tuples $(A_{j},B_{j},q_{j}),\penalty 10000\ 0\leqslant j\leqslant k-1$ for each node $\mathbf{N}_{j},\penalty 10000\ 0\leqslant j\leqslant k-1$ on that path, from the root to $\mathbf{M}$. The quantities $A_{j}$ and $B_{j}$ represent the maximum possible lengths of the two edges emanating from a node $\mathbf{N}_{j}$, and $q$ represents the one-step splitting probability leading to $\mathbf{N}_{j}$, based on the constraints from the separated tree property.
We start from $(A_{0},B_{0},q_{0})=(1,1,1)$. Again, denote the lengths of the edges connecting the node $\mathbf{N}_{j-1}$ to its two descendants at level $j$ by $a_{j},b_{j},\,j\in[k-1]$, where $b_{j}=|\ell_{\mathbf{N}_{j-1}}-\ell_{\mathbf{N}_{j}}|$. Consider the node $\mathbf{N}_{j}$ at level $j$. Then $(A_{j},B_{j},q_{j})$ must be of the form (while not distinguishing the order of $A_{j}$ and $B_{j}$)
@@
$A_{k-1}B_{k-1}\prod_{\ell=0}^{k-1}q_{\ell}=\prod_{\ell=1}^{k-1}\frac{A_{\ell}B_{\ell}}{A_{\ell-1}B_{\ell-1}}\prod_{\ell=1}^{k-1}q_{\ell}\leqslant\prod_{\ell=1}^{k-1}\Big(\frac{1}{q_{\ell}}-1\Big)q_{\ell}=\prod_{\ell=1}^{k-1}(1-q_{\ell}),$
@@
By the martingale property and the main result of Simons (1970),
@@
At level $k$, there are $2^{k}$ vertices (denoted by $j\in[2^{k}]$) in the binary tree representation of the partition-based martingale approximation. Denote the probabilities of the vertices by $p_{j}$ (which correspond to $\mathbb{P}(U\in A_{j}),\penalty 10000\ A_{j}\in\pi_{k}$) and the locations by $\ell_{j}$. It follows that
@@
Note that $\sum_{j}p_{j}=1$ and $\sum_{j}d_{j}\leqslant\sup U-\inf U\leqslant 1$. It follows analogously to the Hölder’s inequality argument in (69) that
@@ B.1 Proofs of results in Section 2
Let $\delta>0$. Then for any joint distribution $(X,Y)$ with $X\geqslant\mathbb{E}[Y]$ or $X\leqslant\mathbb{E}[Y]$, it holds that
@@
By symmetry, we assume that $X\geqslant\mathbb{E}[Y]$. Let $C=(1+\delta)^{3}/\delta^{3}$. Without loss of generality, we assume that $\mathbb{E}[Y]=0$. Equivalent to (43) is
@@
Lemma B.2 (Oscillation growth under cyclic axis-aligned refinement).
Let $\{\pi_{k}\}_{k\geqslant 0}$ be the sequence of partitions generated by the cyclic splitting schedule, and let $k=dq+r$ with $q\in\mathbb{N}_{0}$ and $0\leqslant r<d$. Suppose that
$g(\mathbf{x})=\sum_{i=1}^{d}g_{i}(x_{i})$
is an additive representation with each $g_{i}$ of bounded variation. For a cell $A=\prod_{i=1}^{d}I_{i}(A)\in\pi_{k}$, define
$\mathrm{osc}_{A}(g):=\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x}).$
Then
$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant 2^{\,k-\lfloor k/d\rfloor}\sum_{i=1}^{d}\|g_{i}\|_{\mathrm{TV}}.$
Consequently, after taking the infimum over additive representations of $g$,
$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant 2^{\,k-\lfloor k/d\rfloor}\|g\|_{\mathrm{TV}}.$
Proof.
Define
$\mathrm{osc}_{I}(g_{i}):=\sup_{u\in I}g_{i}(u)-\inf_{u\in I}g_{i}(u).$
For every cell $A=\prod_{i=1}^{d}I_{i}(A)$, additivity gives
$\mathrm{osc}_{A}(g)\leqslant\sum_{i=1}^{d}\mathrm{osc}_{I_{i}(A)}(g_{i}).$
Hence it suffices to control, for each coordinate $i\in[d]$,
$S_{i}(k):=\sum_{A\in\pi_{k}}\mathrm{osc}_{I_{i}(A)}(g_{i}).$
At level $k$, the cyclic rule splits along coordinate
$j_{k}:=1+(k\bmod d).$
If a cell $A$ is not split, then its contribution to every $S_{i}$ is unchanged. If a cell $A$ is split into $A^{L}$ and $A^{R}$ along coordinate $j_{k}$, then for every $i\neq j_{k}$,
$I_{i}(A^{L})=I_{i}(A^{R})=I_{i}(A),$
so the contribution to $S_{i}$ is duplicated:
$\mathrm{osc}_{I_{i}(A^{L})}(g_{i})+\mathrm{osc}_{I_{i}(A^{R})}(g_{i})=2\,\mathrm{osc}_{I_{i}(A)}(g_{i}).$
For the split coordinate $j_{k}$, the child intervals partition the parent interval, hence
$\mathrm{osc}_{I_{j_{k}}(A^{L})}(g_{j_{k}})+\mathrm{osc}_{I_{j_{k}}(A^{R})}(g_{j_{k}})\leqslant\mathrm{osc}_{I_{j_{k}}(A)}(g_{j_{k}}).$
Therefore
$S_{j_{k}}(k+1)\leqslant S_{j_{k}}(k),\qquad S_{i}(k+1)\leqslant 2S_{i}(k)\quad\text{for }i\neq j_{k}.$
Fix $i\in[d]$. By time $k=dq+r$, coordinate $i$ has been selected exactly $q+\mathbf{1}_{\{i\leqslant r\}}$ times, and has not been selected exactly
$k-q-\mathbf{1}_{\{i\leqslant r\}}$
times. Since $S_{i}$ never increases when coordinate $i$ is selected and can at most double otherwise, we obtain
$S_{i}(k)\leqslant 2^{k-q-\mathbf{1}_{\{i\leqslant r\}}}S_{i}(0)\leqslant 2^{k-q}\|g_{i}\|_{\mathrm{TV}},$
because $\mathrm{osc}_{I_{i}(A_{0})}(g_{i})\leqslant\|g_{i}\|_{\mathrm{TV}}$ at level 0.
Summing over $i$ yields
$\sum_{A\in\pi_{k}}\mathrm{osc}_{A}(g)\leqslant\sum_{i=1}^{d}S_{i}(k)\leqslant 2^{\,k-\lfloor k/d\rfloor}\sum_{i=1}^{d}\|g_{i}\|_{\mathrm{TV}}.$
Taking the infimum over all additive representations of $g$ gives the final claim. ∎
@@
To this end, we need an upper bound for $\mathrm{Var}(Y)$. Fix $g\in\mathcal{G}$ and let $g^{*}:=(\sup g+\inf g)/2$. By Minkowski’s inequality,
$\begin{split}\mathrm{Var}(Y)^{1/2}=\min_{y\in\mathbb{R}}\mathbb{E}[(Y-y)^{2}]^{1/2}\leqslant\mathbb{E}[(Y-g^{*})^{2}]^{1/2}&\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\mathbb{E}[(g(\mathbf{X})-g^{*})^{2}]^{1/2}\\ &\leqslant\mathbb{E}[(Y-g(\mathbf{X}))^{2}]^{1/2}+\frac{1}{2}\left\lVert g\right\rVert_{\mathrm{TV}}.\end{split}$ (46)
@@
It remains to establish (45). Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ by the cyclic minimax construction. Note that by construction, $J_{k}\leqslant 2^{k}$ and $M_{k}\mid\mathbf{X}\in I_{j}$ is a piecewise constant random variable, which takes a constant value $y_{k,j}$ inside the box $I_{j}$, so $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$. By definition, $y_{k,j}=\mathbb{E}[Y\mid\mathbf{X}\in I_{j}]$. Define $g_{j}=g|_{I_{j}}$ and $\mathrm{Ran}(g_{j})=[\inf g_{j},\sup g_{j}]$, and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
@@
[…]
@@
[…]
@@
as desired. On the other hand, suppose that $k=dq+r$ for some integer $q$ and $0\leqslant r<d$, and denote by $\pi_{k}$ the resulting partition of $\mathbb{R}^{d}$. We use Lemma B.2 and observe that
@@
Recall that $\Delta Y:=\sup\operatorname{supp}Y-\inf\operatorname{supp}Y$. The only difference from the proof of Theorem 3 is (52), since for non-marginally atomless measures (10) and (11) may not hold. Instead, by our assumptions, (10) is replaced by
$\max\Big\{\varphi_{L}(\hat{x}_{j}),\varphi_{R}(\hat{x}_{j})\Big\}\leqslant\frac{1}{2}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]+\frac{(\Delta Y)^{2}}{N}.$ (56)
To see this, let $B$ and $C$ be disjoint events with $\mathbb{P}(C)=m$, and suppose that $Y\in[a,b]$ almost surely, so $\Delta Y=b-a$. We let $\mu_{B}:=\mathbb{E}[Y\mid B]$. Since the conditional mean minimizes squared loss,
$\mathbb{P}(B\cup C)\mathrm{Var}(Y\mid B\cup C)$ $\leqslant\mathbb{E}[(Y-\mu_{B})^{2}{\mathbbm{1}}_{B\cup C}]$
$=\mathbb{P}(B)\mathrm{Var}(Y\mid B)+\mathbb{E}[(Y-\mu_{B})^{2}{\mathbbm{1}}_{C}]\leqslant\mathbb{P}(B)\mathrm{Var}(Y\mid B)+m(\Delta Y)^{2}.$
Therefore, adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $(\Delta Y)^{2}/N$.
As a consequence of (56), if we denote by
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{(\Delta Y)^{2}}{N},\penalty 10000\ k\geqslant 0;\quad u_{0}=\mathrm{Var}(Y).$